OpenStudy (anonymous):

Complete the following sets of Pythagorean triples. (3, x, 5) (x, 12, 13) (7, x, 25) (8, 15, x) (21, x, 29)

OpenStudy (anonymous):

Just use \(a^2+b^2=c^2\) for each set \((a,b,c)\). For example, the first one, \(a=3, b=x, c=5, 3^2+x^2=5^2, 9+x^2=25, x^2=16, x=4\).